# Quantum teleportation implies symmetry-protected topological order

### v2

@@ 5.2 Single-qubit case: $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$
The proof of Lemma 16 appears in App. N. The Lemma shows that the resource state $\lvert\Psi_{t}\rangle$ (36) has nontrivial string order captured by a set of order parameters (57) corresponding to intervals $\mathcal{I}\subseteq[i,f]$. Valid endpoints are labelled $s=[0,m+1]$, where $s=0$ corresponds to $j\approx i+vT$, $s=m+1$ to $j=f$, and all others to the $m$ nonoverlapping “measurement regions.”
The string order parameters (57) are illustrated in Fig. 5(b). The operator $\Sigma^{\nu}_{1}$ is guaranteed to have no overlap in support with $\mathcal{U}^{\dagger}\bar{A}_{\nu,2}$—the observable immediately to the right of the left endpoint operator in Fig. 5(b)—by Lemma 16. This can be ensured by multiplying the “naïve” choice of $\Sigma^{\nu}_{1}$ by an initial-state stabilizer $\mathsf{S}_{\nu,1}$ to remove any support in triangle defined by the parallelogram and dashed line. The same holds for the right endpoint operator. We also point out that $\Sigma^{\nu}_{1}$ has no overlap with $\sigma^{\nu}_{i}$, and that this extends to any choice of interval $\mathcal{I}=[a,b]$, where the endpoint operators of the corresponding string order parameter $\mathcal{S}_{\nu,a,b}$ (57) have no overlap with those of $\mathcal{S}_{\nu,a\pm 1,b}$ nor $\mathcal{S}_{\nu,a,b\pm 1}$ by Lemma 16.
The string order parameters (57) indicate that $\lvert\Psi_{t}\rangle$ (36) is symmetric (50) with respect to a group $\mathcal{G}=\mathbb{Z}_{2}\times\mathbb{Z}_{2}$, which is projectively represented by the Pauli matrices $\{\mathds{1},X,Y,Z\}$, so that $\Omega(x,z)=-1$ (52). In the bulk, $\mathcal{G}$ has a linear (unitary) representation
$U_{\mathcal{I}}(\nu)=\prod\limits_{s\in\mathcal{I}}\,\bar{A}_{\nu,s}\,,~{}~{}$ (61)
@@
The resource state $\lvert\Psi_{t}\rangle$ (36) can be embedded in a thermodynamically large (i.e., infinite) chain, and the string order (60) extended to arbitrarily large intervals by implementing the truncation procedure of Lemma 16 in reverse.
Corollary 17 follows from Lemma 16 and its proof in App. N. The Corollary establishes that one can use the resource state $\lvert\Psi_{t}\rangle$ (36) to define a family of states of unbounded length, each of which has perfect string order, meaning that the string order survives into the thermodynamic limit. This also establishes that $\lvert\Psi_{t}\rangle$ (36) cannot be prepared in finite time $t\ll d(i,f)$ under local unitary time evolution (with $v=\mathsf{O}\left(1\right)$ finite) that respects the symmetry $\mathcal{G}=\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ Huang and Chen (2015). Thus, the resource state (36) for any $k=1$ standard teleportation protocol $\mathcal{W}$ belongs to a nontrivial SPT phase indicated by $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ string order, where the measured operators $A_{j}$ used to determine $\mathcal{R}_{x/z}$ (7) are those that make up the bulk of the $x/z$ string order parameters (57).
@@ 5.3 Multiqubit case: $\mathbb{Z}_{2}^{2k}$
Consider a protocol with canonical form $\mathcal{W}=\mathcal{R}\,\mathcal{M}\,\mathcal{U}$ (31) that achieves standard teleportation of $k\geq 1$ qubits (see Def. 6). Omitting any unnecessary measurements from $\mathcal{M}$ (see Cor. 14), we identify $m$ measurement regions—where $\bar{A}_{n,\nu,s}$ is a product of all measured observables $\bar{A}_{j}$ (4) in region $s$ attached to $\sigma^{\nu}_{f_{n}}$ (40)—such that $\bar{A}_{n,\nu,s+1}$ has no overlap with the endpoint operator
$\Sigma^{\nu}_{n,s}$ $=\sigma^{\nu}_{i_{n}}\,\mathcal{U}^{\dagger}\,\bar{A}_{n,\nu,1}\cdots\bar{A}_{n,\nu,s}\,\mathcal{U}$
$=\Sigma^{\nu}_{n,s-1}\,\mathcal{U}^{\dagger}\,\bar{A}_{n,\nu,s}\,\mathcal{U}\,,~{}~{}$ (62)
@@
$\mathcal{S}^{n,\nu}_{a,b}\equiv\mathcal{U}\,\Sigma^{\nu}_{n,a}\,\mathcal{U}^{\dagger}\,\bar{A}_{n,\nu,a+1}\cdots\bar{A}_{n,\nu,b}\,\mathcal{U}\,\Sigma^{\nu}_{n,b}\,\mathcal{U}^{\dagger}\,,$ (63)
@@
*$\left\langle*\middle|\Psi_{t}\middle|\,\mathcal{S}^{n,\nu}_{a,b}\,\right\rangle{\Psi_{t}}=1\,,~{}~{}$ (64)
for the resource state $\lvert\Psi_{t}\rangle$ (36), indicating nontrivial string order (see Def. 8) with respect to the symmetry
$\mathcal{G}_{k}=\left(\mathbb{Z}_{2}\times\mathbb{Z}_{2}\right)^{k}=\mathbb{Z}^{2k}_{2}\,,~{}~{}$ (65)
so that $\mathcal{U}$ prepares a resource state $\lvert\Psi_{t}\rangle$ (36) with nontrivial SPT order, $\mathcal{M}$ measures its end-to-end (15) string order parameters $\mathcal{S}^{n,\nu}_{0,m+1}$ (63), and $\mathcal{R}$ corrects the errors $\chi$, which belong to a projective representation of $\mathcal{G}_{k}$ (65).
@@
The case for $k=1$ is proven in Lemmas 15 and 16. By definition, the logical operators for different logical qubits commute; by Cor. 14, all necessary measured observables commute for any $k$, while all other measurements have been omitted without affecting $\mathcal{W}$. Thus, Lemmas 15 and 16 hold for each logical qubit individually (i.e., independently): The preceding arguments guarantee that the string order parameters (63) for distinct logical qubits locally commute. Moreover, Cor. 17 extends to $k>1$ without caveat. As a result, $\lvert\Psi_{t}\rangle$ (36) belongs to a nontrivial SPT phase with $\mathbb{Z}^{2k}_{2}$ string order. ∎
Theorem 18 is the main result of this paper. In the optimal case, the $M=2km$ measurements are applied to the $2k$ sublattices of the chain at the endpoints of each valid subinterval $\mathcal{I}$ described in Sec. 5.1. The same resource state $\lvert\Psi_{t}\rangle$ (36) can be used to teleport logical states $\lvert\psi_{n}\rangle$ between the endpoints of any such interval $\mathcal{I}$; by Cor. 17, $\lvert\Psi_{t}\rangle$ can also be embedded in a thermodynamically large chain with nontrivial string order on infinitely many such endpoints. Hence, the resource state $\lvert\Psi_{t}\rangle$ (36) belongs to a nontrivial SPT phase protected by the symmetry $\mathcal{G}_{k}$ (65), where logical states are teleported between the edge modes of open intervals on the chain by measuring the string order parameters (63) on the intervals $[i_{n},f_{n})$.
This also holds even for suboptimal protocols that realize standard teleportation. In particular, while using more measurements are made than are needed to achieve teleportation distance $L$ (13) complicates the delineation of valid intervals (i.e., endpoints), they are guaranteed to exist by Theorem 18 (and Lemmas 15 and 16). Hence, even suboptimal standard teleportation protocols $\mathcal{W}$ are equivalent to preparing an SPT resource state $\lvert\Psi_{t}\rangle$ (36) protected by $\mathcal{G}=\mathbb{Z}_{2}^{2k}$ and measuring its $2\,k$ string order parameters (63) on the $k$ intervals $[i_{n},f_{n})$.
@@ 6 Example teleportation protocols
We now describe several examples that illustrate our main result. In Sec. 6.1, we consider a $k=1$ teleportation protocol whose resource state is the same cluster state that appears in Sec. 2.4, and we explain how $Y$-basis measurements are nonetheless compatible with Theorem 18. In Sec. 6.2 we explain how hypergraph states Rossi et al. (2013); Gachechiladze et al. (2019); Miller and Miyake (2016) are a viable resource for $k=1$ teleportation, despite not generally being recognized as SPTs and being nonstabilizer states Gottesman (1997). In Sec. 6.3, we generalize the cluster states of Secs. 2.4 and 6.1 to $k\geq 1$, and describe the protecting symmetries and string order parameters. Finally, in Sec. 6.4, we present a family of optimal resource states for standard teleportation of $k\geq 1$ qubits and discuss the reelvant symmetries and order parameters.
Figure 6: A hypergraph lattice with qubits assigned to the $N=17$ vertices. The hypergraph state $\lvert\Psi_{t}\rangle$ recovers from $\lvert\Psi_{0}\rangle=\lvert+++\cdots\rangle$ (68), where $X_{j}\lvert+\rangle=\lvert+\rangle$, by applying CCZ gates to every unshaded triangular face via $\mathcal{U}$ (69).
@@ 6.1 Alternative measurements on the cluster state
We first consider an extension of the cluster state example of Sec. 2.4 Briegel and Raussendorf (2001). However, instead of measuring in the $X$ basis (as in Fig. 2), one can instead measure in the $Y$ basis Stephen et al. (2022). Theorem 18 then implies that the cluster state has nontrivial string order with respect to a unitary symmetry (49) that is diagonal in the on-site $Y$ basis. We now show that this is, indeed, the case, and identify an alternative measurement pattern for $Y$-basis measurements compatible with cluster-state teleportation.
The protocol with $Y$-basis measurements is similar to the $X$-basis protocol in Sec. 2.4. In fact, both protocols have the same physical unitary channel $\mathcal{U}$ (21). To understand the measurement and recovery channels, it is helpful to consider the Heisenberg evolution of the logical operators. In analogy to Eq. 2.4, we require that the logical operators acting on the cluster state $\lvert\Psi_{t}\rangle$ (36) are
$\mathcal{M}^{\dagger}\,\mathcal{R}^{\dagger}\,X_{f}\,\mathcal{R}\,\mathcal{M}$ $=YY\mathds{1}\,YY\mathds{1}\,\cdots YY\mathds{1}\,X_{f}$ (66a)
$\mathcal{M}^{\dagger}\,\mathcal{R}^{\dagger}\,Z_{f}\,\mathcal{R}\,\mathcal{M}$ $=\mathds{1}\,YY\mathds{1}\,YY\cdots\mathds{1}\,YY\,Z_{f}\,,~{}$ (66b)
for chains of length $N=3m+1$ (though it is possible to modify the protocol for other $N$), where we have projected onto the default initial Stinespring state $\lvert\boldsymbol{0}\rangle$. Essentially, one measures all sites except $f$ in the $Y$ basis, where “measurement regions” involve three neighboring sites.
Following Sec. 2.4, we again identify the stabilizers $\mathsf{S}_{x}$ and $\mathsf{S}_{z}$ (2.4) using Eq. 1 of Prop. 1 to ensure successful teleportation. Then Eq. 6.1 implies that
$\mathcal{U}\,\mathsf{S}_{x}\,\mathcal{U}^{\dagger}$ $=Z_{1}\,X_{2}\,(\mathds{1}\,YY\mathds{1}\,YY\cdots\mathds{1}\,)\,X_{N}$ (67a)
$\mathcal{U}\,\mathsf{S}_{z}\,\mathcal{U}^{\dagger}$ $=Z_{1}\,Y_{2}(Y\mathds{1}\,YY\mathds{1}\,Y\cdots Y)\,Z_{N}\,,~{}$ (67b)
@@
Moreover, the operators $\mathsf{S}_{x}$ and $\mathsf{S}_{z}$ (6.1) realize end-to-end string order parameters (15). These two operators commute in the bulk but have anticommuting endpoints. Since they have expectation value one in the cluster state, they imply that the cluster state also has nontrivial string order with respect to the symmetry $\mathcal{G}=\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ whose (bulk) generators are $Y\mathds{1}\,YY\mathds{1}\,Y\cdots$ and $\mathds{1}\,YY\mathds{1}\,YY\cdots$.
Hence, we identify measurement regions as having size three, since this is the minimal translation-invariant unit cell of the symmetry generators. Finally, we note that the number of local measurements is minimized if one measures $Y_{1}Y_{2}$ and $Y_{2}Y_{3}$ in each region; however, one can instead measure only $Y_{1}$ and $Y_{3}$, at the cost of applying an additional FDQC after $\mathcal{U}$ (21). These single-qubit measurements are compatible with Def. 3, and more similar to the $X$-basis protocol discussed in Sec. 2.4, but the measurement regions still have size three.

### v3

@@ 5.2 Single-qubit case: $\mathbb{Z}^{\,}_{2}\times\mathbb{Z}^{\,}_{2}$
The proof of Lemma 16 appears in App. N. The Lemma shows that the resource state $\lvert\Psi^{\,}_{t}\rangle$ (36) has nontrivial string order captured by a set of order parameters (57) corresponding to intervals $\mathcal{I}\subseteq[i,f]$. Valid endpoints are labelled $s=[0,m+1]$, where $s=0$ corresponds to $j\approx i+vT$, $s=m+1$ to $j=f$, and all others to the $m$ nonoverlapping “measurement regions.”
The string order parameters (57) are illustrated in Fig. 5(b). The operator $\Sigma^{\nu}_{1}$ is guaranteed to have no overlap in support with $\mathcal{U}^{\dagger}\bar{A}^{\,}_{\nu,2}$—the observable immediately to the right of the left endpoint operator in Fig. 5(b)—by Lemma 16. This can be ensured by multiplying the “naïve” choice of $\Sigma^{\nu}_{1}$ by an initial-state stabilizer $\mathsf{S}^{\,}_{\nu,1}$ to remove any support in triangle defined by the parallelogram and dashed line. The same holds for the right endpoint operator. We also point out that $\Sigma^{\nu}_{1}$ has no overlap with $\sigma^{\nu}_{i}$, and that this extends to any choice of interval $\mathcal{I}=[a,b]$, where the endpoint operators of the corresponding string order parameter $\mathcal{S}^{\,}_{\nu,a,b}$ (57) have no overlap with those of $\mathcal{S}^{\,}_{\nu,a\pm 1,b}$ nor $\mathcal{S}^{\,}_{\nu,a,b\pm 1}$ by Lemma 16.
The string order parameters (57) indicate that $\lvert\Psi^{\,}_{t}\rangle$ (36) is symmetric (50) with respect to a group $\mathcal{G}=\mathbb{Z}^{\,}_{2}\times\mathbb{Z}^{\,}_{2}$, which is projectively represented by the Pauli matrices $\{\mathds{1},X^{\vphantom{*}},Y^{\vphantom{*}},Z^{\vphantom{*}}\}$, so that $\Omega(x,z)=-1$ (52). In the bulk, $\mathcal{G}$ has a linear (unitary) representation
$U_{\mathcal{I}}(\nu)=\prod\limits_{s\in\mathcal{I}}\,\bar{A}^{\vphantom{\prime}}_{\nu,s}\,,~{}~{}$ (61)
@@
The resource state $\lvert\Psi^{\,}_{t}\rangle$ (36) can be embedded in a thermodynamically large (i.e., infinite) chain, and the string order (60) extended to arbitrarily large intervals by implementing the truncation procedure of Lemma 16 in reverse.
Corollary 17 follows from Lemma 16 and its proof in App. N. The Corollary establishes that one can use the resource state $\lvert\Psi^{\,}_{t}\rangle$ (36) to define a family of states of unbounded length, each of which has perfect string order, meaning that the string order survives into the thermodynamic limit. This also establishes that $\lvert\Psi^{\,}_{t}\rangle$ (36) cannot be prepared in finite time $t\ll d(i,f)$ under local unitary time evolution (with $v=\mathsf{O}\left(1\right)$ finite) that respects the symmetry $\mathcal{G}=\mathbb{Z}^{\,}_{2}\times\mathbb{Z}^{\,}_{2}$ [70]. Thus, the resource state (36) for any $k=1$ standard teleportation protocol $\mathcal{W}$ belongs to a nontrivial SPT phase indicated by $\mathbb{Z}^{\,}_{2}\times\mathbb{Z}^{\,}_{2}$ string order, where the measured operators $A_{j}$ used to determine $\mathcal{R}_{x/z}$ (38) are those that make up the bulk of the $x/z$ string order parameters (57).
@@ 5.3 Multiqubit case: $\mathbb{Z}_{2}^{2k}$
Consider a protocol with canonical form $\mathcal{W}=\mathcal{R}\,\mathcal{M}\,\mathcal{U}$ (31) that achieves standard teleportation of $k\geq 1$ qubits (see Def. 6). Omitting any unnecessary measurements from $\mathcal{M}$ (see Cor. 14), we identify $m$ measurement regions—where $\bar{A}^{\,}_{n,\nu,s}$ is a product of all measured observables $\bar{A}^{\,}_{j}$ (33) in region $s$ attached to $\sigma^{\nu}_{f_{n}}$ (40)—such that $\bar{A}^{\,}_{n,\nu,s+1}$ has no overlap with the endpoint operator
$\Sigma^{\nu}_{n,s}$ $=\sigma^{\nu}_{i_{n}}\,\mathcal{U}^{\dagger}\,\bar{A}^{\vphantom{\prime}}_{n,\nu,1}\cdots\bar{A}^{\vphantom{\prime}}_{n,\nu,s}\,\mathcal{U}$
$=\Sigma^{\nu}_{n,s-1}\,\mathcal{U}^{\dagger}\,\bar{A}^{\vphantom{\prime}}_{n,\nu,s}\,\mathcal{U}\,,~{}~{}$ (62)
@@
$\mathcal{S}^{n,\nu}_{a,b}\equiv\mathcal{U}\,\Sigma^{\nu}_{n,a}\,\mathcal{U}^{\dagger}\,\bar{A}^{\vphantom{\prime}}_{n,\nu,a+1}\cdots\bar{A}^{\vphantom{\prime}}_{n,\nu,b}\,\mathcal{U}\,\Sigma^{\nu}_{n,b}\,\mathcal{U}^{\dagger}\,,$ (63)
@@
$\left\langle*\middle|\Psi^{\vphantom{\prime}}_{t}\middle|\,\mathcal{S}^{n,\nu}_{a,b}\,\right\rangle{\Psi^{\vphantom{\prime}}_{t}}=1\,,~{}~{}$ (64)
for the resource state $\lvert\Psi^{\,}_{t}\rangle$ (36), indicating nontrivial string order (see Def. 8) with respect to the symmetry
$\mathcal{G}_{k}=\left(\mathbb{Z}^{\vphantom{\prime}}_{2}\times\mathbb{Z}^{\vphantom{\prime}}_{2}\right)^{k}=\mathbb{Z}^{2k}_{2}\,,~{}~{}$ (65)
so that $\mathcal{U}$ prepares a resource state $\lvert\Psi^{\,}_{t}\rangle$ (36) with nontrivial SPT order, $\mathcal{M}$ measures its end-to-end (54) string order parameters $\mathcal{S}^{n,\nu}_{0,m+1}$ (63), and $\mathcal{R}$ corrects the errors $\chi$, which belong to a projective representation of $\mathcal{G}_{k}$ (65).
@@
The case for $k=1$ is proven in Lemmas 15 and 16. By definition, the logical operators for different logical qubits commute; by Cor. 14, all necessary measured observables commute for any $k$, while all other measurements have been omitted without affecting $\mathcal{W}$. Thus, Lemmas 15 and 16 hold for each logical qubit individually (i.e., independently): The preceding arguments guarantee that the string order parameters (63) for distinct logical qubits locally commute. Moreover, Cor. 17 extends to $k>1$ without caveat. As a result, $\lvert\Psi^{\,}_{t}\rangle$ (36) belongs to a nontrivial SPT phase with $\mathbb{Z}^{2k}_{2}$ string order. ∎
Theorem 18 is the main result of this paper. In the optimal case, the $M=2km$ measurements are applied to the $2k$ sublattices of the chain at the endpoints of each valid subinterval $\mathcal{I}$ described in Sec. 5.1. The same resource state $\lvert\Psi^{\,}_{t}\rangle$ (36) can be used to teleport logical states $\lvert\psi_{n}\rangle$ between the endpoints of any such interval $\mathcal{I}$; by Cor. 17, $\lvert\Psi^{\,}_{t}\rangle$ can also be embedded in a thermodynamically large chain with nontrivial string order on infinitely many such endpoints. Hence, the resource state $\lvert\Psi^{\,}_{t}\rangle$ (36) belongs to a nontrivial SPT phase protected by the symmetry $\mathcal{G}_{k}$ (65), where logical states are teleported between the edge modes of open intervals on the chain by measuring the string order parameters (63) on the intervals $[i_{n},f_{n})$.
This also holds even for suboptimal protocols that realize standard teleportation. In particular, while using more measurements are made than are needed to achieve teleportation distance $L$ (13) complicates the delineation of valid intervals (i.e., endpoints), they are guaranteed to exist by Theorem 18 (and Lemmas 15 and 16). Hence, even suboptimal standard teleportation protocols $\mathcal{W}$ are equivalent to preparing an SPT resource state $\lvert\Psi^{\,}_{t}\rangle$ (36) protected by $\mathcal{G}=\mathbb{Z}_{2}^{2k}$ and measuring its $2\,k$ string order parameters (63) on the $k$ intervals $[i_{n},f_{n})$.
@@ 6 Example teleportation protocols
We now describe several examples that illustrate our main result. In Sec. 6.1, we consider a $k=1$ teleportation protocol whose resource state is the same cluster state that appears in Sec. 2.4, and we explain how $Y^{\vphantom{*}}$-basis measurements are nonetheless compatible with Theorem 18. In Sec. 6.2 we explain how hypergraph states [52, 53, 25] are a viable resource for $k=1$ teleportation, despite not generally being recognized as SPTs and being nonstabilizer states [51]. In Sec. 6.3, we generalize the cluster states of Secs. 2.4 and 6.1 to $k\geq 1$, and describe the protecting symmetries and string order parameters. Finally, in Sec. 6.4, we present a family of optimal resource states for standard teleportation of $k\geq 1$ qubits and discuss the reelvant symmetries and order parameters.
Figure 6: A hypergraph lattice with qubits assigned to the $N=17$ vertices. The hypergraph state $\lvert\Psi^{\,}_{t}\rangle$ recovers from $\lvert\Psi^{\,}_{0}\rangle=\lvert+++\cdots\rangle$ (68), where $X_{j}\lvert+\rangle=\lvert+\rangle$, by applying CCZ gates to every unshaded triangular face via $\mathcal{U}$ (69).
@@ 6.1 Alternative measurements on the cluster state
We first consider an extension of the cluster state example of Sec. 2.4 [43]. However, instead of measuring in the $X^{\vphantom{*}}$ basis (as in Fig. 2), one can instead measure in the $Y^{\vphantom{*}}$ basis [72]. Theorem 18 then implies that the cluster state has nontrivial string order with respect to a unitary symmetry (49) that is diagonal in the on-site $Y^{\vphantom{*}}$ basis. We now show that this is, indeed, the case, and identify an alternative measurement pattern for $Y^{\vphantom{*}}$-basis measurements compatible with cluster-state teleportation.
The protocol with $Y^{\vphantom{*}}$-basis measurements is similar to the $X^{\vphantom{*}}$-basis protocol in Sec. 2.4. In fact, both protocols have the same physical unitary channel $\mathcal{U}$ (21). To understand the measurement and recovery channels, it is helpful to consider the Heisenberg evolution of the logical operators. In analogy to Eq. 25, we require that the logical operators acting on the cluster state $\lvert\Psi^{\,}_{t}\rangle$ (36) are
$\mathcal{M}^{\dagger}\,\mathcal{R}^{\dagger}\,X^{\vphantom{*}}_{f}\,\mathcal{R}\,\mathcal{M}$ $=YY\mathds{1}\,YY\mathds{1}\,\cdots YY\mathds{1}\,X^{\vphantom{*}}_{f}$ (66a)
$\mathcal{M}^{\dagger}\,\mathcal{R}^{\dagger}\,Z^{\vphantom{*}}_{f}\,\mathcal{R}\,\mathcal{M}$ $=\mathds{1}\,YY\mathds{1}\,YY\cdots\mathds{1}\,YY\,Z^{\vphantom{*}}_{f}\,,~{}$ (66b)
for chains of length $N=3m+1$ (though it is possible to modify the protocol for other $N$), where we have projected onto the default initial Stinespring state $\lvert\boldsymbol{0}\rangle$. Essentially, one measures all sites except $f$ in the $Y^{\vphantom{*}}$ basis, where “measurement regions” involve three neighboring sites.
Following Sec. 2.4, we again identify the stabilizers $\mathsf{S}^{\,}_{x}$ and $\mathsf{S}^{\,}_{z}$ (27) using Eq. 8 of Prop. 1 to ensure successful teleportation. Then Eq. 66 implies that
$\mathcal{U}\,\mathsf{S}^{\vphantom{\prime}}_{x}\,\mathcal{U}^{\dagger}$ $=Z^{\vphantom{*}}_{1}\,X^{\vphantom{*}}_{2}\,(\mathds{1}\,YY\mathds{1}\,YY\cdots\mathds{1}\,)\,X^{\vphantom{*}}_{N}$ (67a)
$\mathcal{U}\,\mathsf{S}^{\vphantom{\prime}}_{z}\,\mathcal{U}^{\dagger}$ $=Z^{\vphantom{*}}_{1}\,Y^{\vphantom{*}}_{2}(Y\mathds{1}\,YY\mathds{1}\,Y\cdots Y)\,Z^{\vphantom{*}}_{N}\,,~{}$ (67b)
@@
Moreover, the operators $\mathsf{S}^{\,}_{x}$ and $\mathsf{S}^{\,}_{z}$ (67) realize end-to-end string order parameters (54). These two operators commute in the bulk but have anticommuting endpoints. Since they have expectation value one in the cluster state, they imply that the cluster state also has nontrivial string order with respect to the symmetry $\mathcal{G}=\mathbb{Z}^{\,}_{2}\times\mathbb{Z}^{\,}_{2}$ whose (bulk) generators are $Y\mathds{1}\,YY\mathds{1}\,Y\cdots$ and $\mathds{1}\,YY\mathds{1}\,YY\cdots$.
Hence, we identify measurement regions as having size three, since this is the minimal translation-invariant unit cell of the symmetry generators. Finally, we note that the number of local measurements is minimized if one measures $Y^{\vphantom{*}}_{1}Y^{\vphantom{*}}_{2}$ and $Y^{\vphantom{*}}_{2}Y^{\vphantom{*}}_{3}$ in each region; however, one can instead measure only $Y^{\vphantom{*}}_{1}$ and $Y^{\vphantom{*}}_{3}$, at the cost of applying an additional FDQC after $\mathcal{U}$ (21). These single-qubit measurements are compatible with Def. 3, and more similar to the $X^{\vphantom{*}}$-basis protocol discussed in Sec. 2.4, but the measurement regions still have size three.